# Multi-Objective Optimization for Sparse Deep Multi-Task Learning

### v3

@@ Abstract
Different conflicting optimization criteria arise naturally in various Deep Learning scenarios. These can address different main tasks (i.e., in the setting of Multi-Task Learning), but also main and secondary tasks such as loss minimization versus sparsity. The usual approach is a simple weighting of the criteria, which formally only works in the convex setting. In this paper, we present a Multi-Objective Optimization algorithm using a modified Weighted Chebyshev scalarization for training Deep Neural Networks (DNNs) with respect to several tasks. By employing this scalarization technique, the algorithm can identify all optimal solutions of the original problem while reducing its complexity to a sequence of single-objective problems. The simplified problems are then solved using an Augmented Lagrangian method, enabling the use of popular optimization techniques such as Adam and Stochastic Gradient Descent, while efficaciously handling constraints. Our work aims to address the (economical and also ecological) sustainability issue of DNN models, with a particular focus on Deep Multi-Task models, which are typically designed with a very large number of weights to perform equally well on multiple tasks. Through experiments conducted on two Machine Learning datasets, we demonstrate the possibility of adaptively sparsifying the model during training without significantly impacting its performance, if we are willing to apply task-specific adaptations to the network weights. The code is available at https://github.com/salomonhotegni/MDMTN
@@ II Related Works
Multi-objective optimization is a technique for optimizing several conflicting objectives simultaneously. A Multi-Objective Optimization Problem (MOP) involves determining the set of optimal compromises when no specific solution can optimize all objectives simultaneously. To gain a deeper understanding of the Multi-objective optimization problems, we recommend [8] and [31]. Most of the MOO algorithms handle optimization problems using approaches like evolutionary algorithms [9, 23]. scalarization methods [6, 14], set-oriented techniques [43, 37], and gradient-based approaches [15]. In this work, we focus on scalarization-based algorithms, that allow for preferences in MOO [6]. Users preferences are integrated by assigning corresponding weights (importance) to each objective. While linear scalarization algorithms struggle to find optimal solutions in non-convex Pareto front cases, in Reinforcement Learning,[46] recommend the Weighted Chebyshev scalarization approach in a Multi-Objective Q-learning framework. This scalarization method uses the Chebyshev distance metric to minimize the distance between the objective functions and a given reference point. [2] solved multi-objective infinite-dimenisonal parameter optimization problems using the Pascoletti-Serafini (PS) scalarization approach. As in our work, the obtained scalar optimization problems are solved using the Augmented Lagrangian method. In [14], it has been proved that this PS scalarization approach is equivalent to the WC scalarization, where a variation of the weights and the reference point in WC correspond respectively to a variation of the direction and the reference point in the PS method. To address the inflexibility of the PS scalarization’s constraints, [1] revised a modified PS scalarization algorithm. Nevertheless, for complex MOO problems, the suggested modified PS scalarization algorithm has yet to be tested. These methods lead to the identification of the Pareto set, where each Pareto optimal point is obtained under a different choice of the scalarization parameter.
@@ III-B Weighted Chebyshev Scalarization
where $\epsilon$ is a sufficiently small positive scalar, constant for all objectives, and can be interpreted as a permissible disturbance of the preference cone $\mathbb{R}_{+}^{m+1}$ [19]. In Appendix A, we present an advanced theorem that guarantees solving problem (Modified WC(k,a)) always leads to Pareto optimal solutions.
@@ III-D Group Ordered Weighted $l_{1}$ (GrOWL)
where $\theta^{(l)}\in\mathbb{R}_{+}^{N_{l-1}}$, $\theta_{1}^{(l)}>0$, and $\theta_{1}^{(l)}\geq\theta_{2}^{(l)}\geq\dotsm\geq\theta_{N_{l-1}}^{(l)}\geq 0.$ We define the objective function $\mathcal{L}_{0}$ as a sum of the layer-wise GrOWL penalties:
@@ III-E Deep Multi-Task Learning Model
A Multi-Task Learning model is required when addressing more than one main task $(m>1)$. In the most frequently used architecture, hard parameter sharing, the hidden layers are completely shared across all tasks, while maintaining multiple task-specific output layers [44]. However, this method is only effective when the tasks are closely associated. Other alternatives have been proposed, including Multi-Task Attention Network (MTAN) [27], Knowledge Distillation for Multi-Task Learning (KDMTL) [26], and Deep Safe Multi-Task Learning (DSMTL) [49].
These approaches are based on the idea of allowing each task to have its fair opportunity to contribute, while simultaneously learning shared features. However, as a soft parameter sharing approach, scalability is an issue with MTAN, since the network’s complexity expands linearly with the number of tasks [47]. On the other hand, KDMTL and DSMTL involve duplicating the shared network for every single task. In this paper, we present a novel Deep MTL model in which the task-specific information that the shared network may miss are learned using task-specific monitors. These monitors are sufficiently simple to ensure that they do not considerably increase the model’s complexity while dealing with many tasks. Our proposed model, namely the Monitored Deep Multi-Task Network (MDMTN), is shown in Figure 1.
@@
7:       GrOWL: Remove insignificant neurons to obtain a model with a sparsity rate $s_{k}$ and save the model if $s_{k}\geq s_{min}$ with better performance on the main tasks.
@@ IV Experiments
We start with the MultiMNIST dataset[42], derived from the MNIST dataset of handwritten digits [25] by gluing two images together, with some overlap. The primary tasks are the classification of the top left and bottom right digits of each image. To study a more complicated problem with less aligned tasks, we have created our own dataset, namely the Cifar10Mnist dataset [^1] See sample images in Appendix A, which contains padded MNIST digits on top of the CIFAR-10 images [24]. On this new dataset, we investigate two primary objectives: identifying the background CIFAR-10 image and classifying the padded MNIST digit of each image. Since we are dealing here with two different data sources (CIFAR-10 and MNIST), that have distinct characteristics and representations, it is challenging to simultaneously learn meaningful features and representations for both tasks. The MultiMNIST dataset, on the other hand, typically involves only one data source, making it less complex in terms of data diversity. These two types of MTL datasets then allows us to generalize our approach more effectively. [44]’s approach (MGDA) and the Hard Parameter Sharing (HPS) model architecture are used as a baseline algorithm and a baseline MTL model. The metrics used to evaluate the model architectures are: Sparsity Rate (SR) = #$\frac{\#zero\_params}{\#total\_params}$, Parameter Sharing (PS) = #$\frac{\#nonzero\_params}{\#unique\_params}$, and Compression Rate (CR) = #$\frac{\#total\_params}{\#unique\_params}$.
@@ IV-A Sparsification Effects on Model Performance
To investigate the impact of the secondary objective (sparsification) on the primary objectives, we examine the 3D Pareto front obtained with all objective functions and its 2D projection onto the primary objectives space by evaluating 90 different preference vectors $k=(k_{0},k_{1},k_{2})$ and a fixed reference point $a=(0,0,0)$. The sparsity coefficient $k_{0}$ values used are $0,10^{-1},10^{-2},10^{-3}$ and $10^{-4}$. Using this approach, we obtain $17$ to $19$ points for each value of $k_{0}$, i.e., $90$ points in total. Due to stochasticity during training, we had to perform a consecutive $\epsilon$-nondominance test [38], to filter out the points that are non-dominated (within a tolerance of $\epsilon$). Appendix A contains the model architectures considered for both datasets and other implementation details. We study the Pareto set and identify the single point with the best overall performance when averaging over all the main tasks. Given the exceptionally high dimensions of MTL models, MOEA/D [51], as an evolutionary algorithm, struggles to yield meaningful results on the primary tasks and is consequently omitted from our report.
@@
  Dataset Architecture Method (SR; CR; PS) Accuracy Training time Task 1 Task 2 (average) (minutes)   MultiMNIST HPS MGDA $(0\%;1;1)$ $97.26\%$ $95.90\%$ $96.580\%$ $4$ Ours: $k=(0.0,0.4,0.6)$ $(0\%;1;1)$ $96.74\%$ $95.53\%$ $96.134\%$ $4$ Ours: $k=(10^{-4},0.5,0.4999)$ $(11.83\%;1.13;1)$ $96.78\%$ $96.14\%$ $96.46\%$ 3 MDMTN (Ours) MGDA $(0\%;1;1)$ $96.88\%$ 96.63% $96.755\%$ $6$ Ours: $k=(0,0.35,0.65)$ $(0\%;1;1)$ $97.10\%$ $96.39\%$ $96.744\%$ $9$ Ours: $k=(10^{-3},0.2,0.799)$ (57.26%;2.7;1.15) 97.29% $96.40\%$ 96.845% $7$ KDMTL $(0\%;1;1)$ $96.58\%$ $95.10\%$ $95.84\%$ $27$ MTAN $(0\%;1;1)$ $96.51\%$ $93.46\%$ $94.985\%$ $29$ STL $(0\%;1;1)$ $97.23\%$ $95.90\%$ $96.565\%$ $19$   Cifar10Mnist HPS MGDA $(0\%;1;1)$ $45.75\%$ $96.78\%$ $71.265\%$ $7$ Ours: $k=(0.0,0.5,0.5)$ $(0\%;1;1)$ $54.00\%$ $95.14\%$ $74.570\%$ $3$ Ours: $k=(10^{-1},0.6,0.3)$ $(13.03;1.25;1.10)$ $56.64\%$ $94.81\%$ $75.724\%$ 2 MDMTN (Ours) MGDA $(0\%;1;1)$ $58.17\%$ 97.04% $77.605\%$ $15$ Ours: $k=(0,0.55,0.45)$ $(0\%;1;1)$ $59.25\%$ $95.19\%$ $77.22\%$ $12$ Ours: $k=(10^{-2},0.8,0.19)$ (29.31%;1.94;1.37) 61.05% $94.26\%$ 77.654% $10$ KDMTL $(0\%;1;1)$ $54.00\%$ $94.35\%$ $74.175\%$ $19$ MTAN $(0\%;1;1)$ $45.30\%$ $93.22\%$ $69.26\%$ $25$ STL $(0\%;1;1)$ $58.11\%$ 97.07% $77.59\%$ $10$
@@
Since the model sparsity tolerance range is small, we do not study a 3D Pareto front for this problem. We show the 2D projection of the obtained Pareto optimal points in Figure 3. Each of the $35$ preference vectors produces a model with a sparsity rate of about $29\%$, and their performances vary between $74.46\%$ and $77.65\%$. When we discard the sparsity objective function, we get models with performance in the range of $75.69\%-77.22\%$. Therefore, when dealing with significantly dissimilar tasks, the effect of sparsity on model performance could display randomness, although it holds the potential for improvement. We intend to further investigate this aspect in our upcoming research endeavors.
@@
Remarkably, our method typically demonstrates shorter training times compared to alternative approaches on both datasets, underscoring its efficiency and computational advantage. Figure 4 shows how efficient a sparse model obtained with GrOWL is, compared to the original one, in generating the Pareto front for the main tasks (see Appendix B for details).
@@
Our MDMTN model architecture excels at finding solutions that simultaneously optimize the main tasks while incorporating sparsity $(k_{0}\neq 0)$. Therefore, MDMTN inherently possesses a stronger ability to capture task-related information while still minimizing unnecessary redundancy in the learned parameters. Furthermore, the fact that both MDMTN Pareto fronts $(k_{0}\neq 0,\ \text{and}\ k_{0}=0)$ outperform those of HPS demonstrates that our model architecture excels at establishing an appropriate balance between primary task performance, regardless of whether sparsity is explicitly considered as an objective.

### v4

@@ Abstract
Different conflicting optimization criteria arise naturally in various Deep Learning scenarios. These can address different main tasks (i.e., in the setting of Multi-Task Learning), but also main and secondary tasks such as loss minimization versus sparsity. The usual approach is a simple weighting of the criteria, which formally only works in the convex setting. In this paper, we present a Multi-Objective Optimization algorithm using a modified Weighted Chebyshev scalarization for training Deep Neural Networks (DNNs) with respect to several tasks. By employing this scalarization technique, the algorithm can identify all optimal solutions of the original problem while reducing its complexity to a sequence of single-objective problems. The simplified problems are then solved using an Augmented Lagrangian method, enabling the use of popular optimization techniques such as Adam and Stochastic Gradient Descent, while efficaciously handling constraints. Our work aims to address the (economical and also ecological) sustainability issue of DNN models, with a particular focus on Deep Multi-Task models, which are typically designed with a very large number of weights to perform equally well on multiple tasks. Through experiments conducted on two Machine Learning datasets, we demonstrate the possibility of adaptively sparsifying the model during training without significantly impacting its performance, if we are willing to apply task-specific adaptations to the network weights. Code is available at https://github.com/salomonhotegni/MDMTN.
@@ II Related Works
Multi-objective optimization is a technique for optimizing several conflicting objectives simultaneously. A Multi-Objective Optimization Problem (MOP) involves determining the set of optimal compromises when no specific solution can optimize all objectives simultaneously. To gain a deeper understanding of the Multi-objective optimization problems, we recommend [8] and [31]. Most of the MOO algorithms handle optimization problems using approaches like evolutionary algorithms [9, 23]. scalarization methods [6, 14], set-oriented techniques [44, 38], and gradient-based approaches [15]. In this work, we focus on scalarization-based algorithms, that allow for preferences in MOO [6]. Users preferences are integrated by assigning corresponding weights (importance) to each objective. While linear scalarization algorithms struggle to find optimal solutions in non-convex Pareto front cases, in Reinforcement Learning,[47] recommend the Weighted Chebyshev scalarization approach in a Multi-Objective Q-learning framework. This scalarization method uses the Chebyshev distance metric to minimize the distance between the objective functions and a given reference point. [2] solved multi-objective infinite-dimenisonal parameter optimization problems using the Pascoletti-Serafini (PS) scalarization approach. As in our work, the obtained scalar optimization problems are solved using the Augmented Lagrangian method. In [14], it has been proved that this PS scalarization approach is equivalent to the WC scalarization, where a variation of the weights and the reference point in WC correspond respectively to a variation of the direction and the reference point in the PS method. To address the inflexibility of the PS scalarization’s constraints, [1] revised a modified PS scalarization algorithm. Nevertheless, for complex MOO problems, the suggested modified PS scalarization algorithm has yet to be tested. These methods lead to the identification of the Pareto set, where each Pareto optimal point is obtained under a different choice of scalarization parameter.
@@ III-B Weighted Chebyshev Scalarization
where $\epsilon$ is a sufficiently small positive scalar, constant for all objectives, and can be interpreted as a permissible disturbance of the preference cone $\mathbb{R}_{+}^{m+1}$ [19]. [20] demonstrated that solving problem (Modified WC(k,a)) always yields Pareto optimal solutions (proper Pareto Optimality).
@@ III-D Group Ordered Weighted $l_{1}$ (GrOWL)
where $\theta^{(l)}\in\mathbb{R}_{+}^{N_{l-1}}$, $\theta_{1}^{(l)}>0$, and $\theta_{1}^{(l)}\geq\theta_{2}^{(l)}\geq\dotsm\geq\theta_{N_{l-1}}^{(l)}\geq 0.$ In this paper, we consider the GrOWL-Spike weight pattern[35], which prioritizes the largest term while weighting the rest equally:
$\begin{cases}\theta_{1}=\beta_{1}+\beta_{2}\\ \theta_{i}=\beta_{2},\ \text{for $i=2,...,N_{l-1}$ }\\ \beta_{1},\ \beta_{2}>0.\end{cases}$ (GrOWL-Spike)
We define the objective function $\mathcal{L}_{0}$ as a sum of the layer-wise GrOWL penalties:
@@ III-E Deep Multi-Task Learning Model
A Multi-Task Learning model is required when addressing more than one main task $(m>1)$. In the most frequently used architecture, hard parameter sharing, the hidden layers are completely shared across all tasks, while maintaining multiple task-specific output layers [45]. However, this method is only effective when the tasks are closely associated. Other alternatives have been proposed, including Multi-Task Attention Network (MTAN) [27], Knowledge Distillation for Multi-Task Learning (KDMTL) [26], and Deep Safe Multi-Task Learning (DSMTL) [50]. These approaches are based on the idea of allowing each task to have its fair opportunity to contribute, while simultaneously learning shared features. However, as a soft parameter sharing approach, scalability is an issue with MTAN, since the network’s complexity expands linearly with the number of tasks [48]. On the other hand, KDMTL and DSMTL involve duplicating the shared network for every single task. In this paper, we present a novel Deep MTL model in which the task-specific information that the shared network may miss are learned using task-specific monitors. These monitors are sufficiently simple to ensure that they do not considerably increase the model’s complexity while dealing with many tasks. Our proposed model, namely the Monitored Deep Multi-Task Network (MDMTN), is shown in Figure 1.
@@
7:       GrOWL: Remove insignificant neurons to obtain a model with a sparsity rate $s_{j}$ and save the model if $s_{j}\geq s_{min}$ with better performance on the main tasks.
@@
Considering a 2D (reshaped) weight matrix $W_{l}$ of the $l^{th}$ layer, the rows represent the outputs from the $(l-1)^{th}$ layer’s neurons. Thus, given a threshold $\tau$, the non-significant rows (in terms of norm) are set to zero at each epoch during the first stage of the training phase, and the model is saved if the average accuracy on the primary tasks improves. For a model to be saved, we define a minimal sparsity rate $s_{min}$. We set a maximum sparsity rate $s_{l_{max}}$ for each layer to avoid irrelevant layers (or zero layers). At the end of this stage, we obtain a sparse model, ignoring the connectivities involving the neurons associated to the zero rows of the weight matrices. For the identification of the clusters of neurons that should be forced to produce identical outputs during the second stage of the training phase, we follow [51] and use the Affinity Propagation method based on the following pairwise similarity metric:
$\mathcal{S}_{l}(i,j)=\dfrac{W_{l,i}^{T}W_{l,j}}{max\big{(}||W_{l,i}||_{2}^{2},||W_{l,j}||_{2}^{2})}\in[-1,1]$
with $W_{l,i}$ and $W_{l,j}$, respectively, the $i^{th}$ and $j^{th}$ rows of $W_{l}$.
In the subsequent phase, the most recent model obtained in the initial stage advances along with the clustering information of its weight matrices. This model undergoes a retraining stage, focusing solely on the interconnections among its important neurons, and at each epoch, for each weight matrix $W_{l}$, the rows belonging to the same cluster are substituted with their mean values. As a result, the final model is a compressed and sparse version of the original one.
@@ IV Experiments
We start with the MultiMNIST dataset[43], derived from the MNIST dataset of handwritten digits [25] by gluing two images together, with some overlap. The primary tasks are the classification of the top left and bottom right digits of each image. To study a more complicated problem with less aligned tasks, we have created our own dataset, namely the Cifar10Mnist dataset (Figure 2), which contains padded MNIST digits on top of the CIFAR-10 images [24].
Figure 2: Sample images from the Cifar10Mnist dataset
On this new dataset, we investigate two primary objectives: identifying the background CIFAR-10 image and classifying the padded MNIST digit of each image. Since we are dealing here with two different data sources (CIFAR-10 and MNIST), that have distinct characteristics and representations, it is challenging to simultaneously learn meaningful features and representations for both tasks. The MultiMNIST dataset, on the other hand, typically involves only one data source, making it less complex in terms of data diversity. These two types of MTL datasets then allows us to generalize our approach more effectively. [45]’s approach (MGDA) and the Hard Parameter Sharing (HPS) model architecture are used as a baseline algorithm and a baseline MTL model. The metrics used to evaluate the model architectures are: Sparsity Rate (SR) = #$\frac{\#zero\_params}{\#total\_params}$, Parameter Sharing (PS) = #$\frac{\#nonzero\_params}{\#unique\_params}$, and Compression Rate (CR) = #$\frac{\#total\_params}{\#unique\_params}$.
@@ IV-A Sparsification Effects on Model Performance
  Dataset Architecture Method (SR; CR; PS) Accuracy Training time Task 1 Task 2 (average) (minutes)   MultiMNIST HPS MGDA $(0\%;1;1)$ $97.26\%$ $95.90\%$ $96.580\%$ $17$ Ours: $k=(0.0,0.4,0.6)$ $(0\%;1;1)$ $96.74\%$ $95.53\%$ $96.134\%$ $5$ Ours: $k=(10^{-4},0.5,0.4999)$ $(11.83\%;1.13;1)$ $96.78\%$ $96.14\%$ $96.46\%$ 4 MDMTN (Ours) MGDA $(0\%;1;1)$ $96.88\%$ 96.63% $96.755\%$ $41$ Ours: $k=(0,0.35,0.65)$ $(0\%;1;1)$ $97.10\%$ $96.39\%$ $96.744\%$ $12$ Ours: $k=(10^{-3},0.2,0.799)$ (57.26%;2.7;1.15) 97.29% $96.40\%$ 96.845% $9$ KDMTL $(0\%;1;1)$ $96.16\%$ $95.56\%$ $95.86\%$ $19$ MTAN $(0\%;1;1)$ $96.40\%$ $95.27\%$ $95.835\%$ $23$ STL $(0\%;1;1)$ $97.23\%$ $95.90\%$ $96.565\%$ $16$   Cifar10Mnist HPS MGDA $(0\%;1;1)$ $45.75\%$ $96.78\%$ $71.265\%$ $9$ Ours: $k=(0.0,0.5,0.5)$ $(0\%;1;1)$ $54.00\%$ $95.14\%$ $74.570\%$ $6$ Ours: $k=(10^{-1},0.6,0.3)$ $(13.03;1.25;1.10)$ $56.64\%$ $94.81\%$ $75.724\%$ 5 MDMTN (Ours) MGDA $(0\%;1;1)$ $58.17\%$ 97.04% $77.605\%$ $23$ Ours: $k=(0,0.55,0.45)$ $(0\%;1;1)$ $59.25\%$ $95.19\%$ $77.22\%$ $22$ Ours: $k=(10^{-2},0.8,0.19)$ (29.31%;1.94;1.37) 61.05% $94.26\%$ 77.654% $19$ KDMTL $(0\%;1;1)$ $58.47\%$ $96.18\%$ $77.325\%$ $23$ MTAN $(0\%;1;1)$ $52.91\%$ $95.58\%$ $74.245\%$ $25$ STL $(0\%;1;1)$ $58.11\%$ 97.07% $77.59\%$ $23$
To investigate the impact of the secondary objective (sparsification) on the primary objectives, we examine the 3D Pareto front obtained with all objective functions and its 2D projection onto the primary objectives space by evaluating 90 different preference vectors $k=(k_{0},k_{1},k_{2})$ and a fixed reference point $a=(0,0,0)$. The sparsity coefficient $k_{0}$ values used are $0,10^{-1},10^{-2},10^{-3}$ and $10^{-4}$. Using this approach, we obtain $17$ to $19$ points for each value of $k_{0}$, i.e., $90$ points in total. Due to stochasticity during training, we had to perform a consecutive $\epsilon$-nondominance test [39], to filter out the points that are non-dominated (within a tolerance of $\epsilon$). The Shared Network module in all MTL architectures used is consistent across each dataset. We study the Pareto set and identify the single point with the best overall performance when averaging over all the main tasks. Given the exceptionally high dimensions of MTL models, MOEA/D [52], as an evolutionary algorithm, struggles to yield meaningful results on the primary tasks and is consequently omitted from our report.
@@
Since the model sparsity tolerance range is small, we do not study a 3D Pareto front for this problem. We show the 2D projection of the obtained Pareto optimal points in Figure 4. Each of the $35$ preference vectors produces a model with a sparsity rate of about $29\%$, and their performances vary between $74.46\%$ and $77.65\%$.
@@
When we discard the sparsity objective function, we get models with performance in the range of $75.69\%-77.22\%$. Therefore, when dealing with significantly dissimilar tasks, the effect of sparsity on model performance could display randomness, although it holds the potential for improvement. We intend to further investigate this aspect in our upcoming research endeavors.
We report the training time for our method’s retraining phase, KDMTL’s main training phase, MGDA training repetitions, MTAN’s training, and the combined training times of STL method for the main tasks. Remarkably, our method typically demonstrates shorter training times compared to alternative approaches on both datasets, underscoring its efficiency and computational advantage.
Figure 5 shows how efficient a sparse model obtained with GrOWL is, compared to the original one, in generating the Pareto front for the main tasks. Our MDMTN model architecture excels at finding solutions that simultaneously optimize the main tasks while incorporating sparsity $(k_{0}\neq 0)$. Therefore, MDMTN inherently possesses a stronger ability to capture task-related information while still minimizing unnecessary redundancy in the learned parameters. Furthermore, the fact that both MDMTN Pareto fronts $(k_{0}\neq 0,\ \text{and}\ k_{0}=0)$ outperform those of HPS demonstrates that our model architecture excels at establishing an appropriate balance between primary task performance, regardless of whether sparsity is explicitly considered as an objective.
@@ V Conclusion
Acknowledgment
This project received funding from the German Federal Ministry of Education and Research (BMBF) through the AI junior research group “Multicriteria Machine Learning”.